# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v3

@@ B.1 Relationships between Ground-truth and Estimated Latent Variables
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})$ denotes all pairwise indice such that $z_{t,i},z_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $4n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ different values of ${\mathbf{z}}_{t-1,m}$, such that the $4n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ values of vector functions $w(m)$ are linearly independent.
@@ B.2 Extension to Multiple Lags and Sequence Lengths
For the sake of simplicity, we consider only one special case with $\tau=1$ and $L=2$ in Theorem A2. Our identifiability theorem can be actually extended to arbitrary lags and subsequences easily. For any given $\tau$, and subsequence which is centered at ${\mathbf{z}}_{t}$ with previous $lo$ and following $hi$ steps, i.e., ${\mathbf{z}}_{t}=\{{\mathbf{z}}_{t-lo},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+hi-1}\}\in\mathbb{R}^{(lo+hi)\times n}$. In this case, the vector function $w(i,j,m)$ in Sufficient Variability Assumption should be modified as
@@ B.3 Identifiability of Latent Variables
First, we demonstrate that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero (zhang2024causal). By contradiction, if the product of the diagonal entry of an invertible matrix $A$ is zero for every row permutation, then we have Equation
@@ Appendix C Discussion of Assumptions
First, The smooth and positive density assumption is standard in the literature on nonlinear ICA (khemakhem2020variational; yao2022temporally; kong2022partial), meaning that the latent variables ${\mathbf{z}}_{t}$ change continuously based on historical information. For instance, in a weather dataset, temperature varies smoothly over time. However, this assumption may be violated if we cannot fully capture the transition probabilities from the observations. To mitigate this, we can sample a larger dataset to better estimate the latent causal process.
Second, the sufficient variability assumption has also been commonly adopted for the existing identifiable results of temporally causal representation learning (yao2021learning; yao2022temporally; chen2024caring). This assumption describes the changeability of latent variables. Take human motion forecasting as an example, the latent variable may represent joint location at different time steps. The linear independence of the latent variables means that the changes of each joint cannot be linearly represented by others. Besides, while the sufficiency assumption is foundational to the theory of identifiability, it is not excessively restrictive. Even in cases where this assumption is not fully satisfied, it is still possible to achieve a degree of subspace identifiability (kong2022partial).
Finally, the sparse latent process is the key assumption of our method, which is common in real-world scenarios. Intuitively, in human motion forecasting, the joints of humans can be considered as latent variables. Since there are few connections among the joints of the human body, the latent process described by the skeleton motion trajectory is sparse. Even though the sparsity assumption is not completely satisfied, we can still attain a subspace level of identifiability (li2024subspace). In this case, each true variable can be a function of, at most, an estimated version of its corresponding variable and those within the intimate set. Let us provide a simple example here. In a video of a moving car, it might be hard to have individual identifiability of the separate car wheels and car body; however, they can be considered as essential parts of the macro variable ’car’. This macro representation might be sufficient for the purpose of modeling the interactions between the car and other objects.
@@ E.3 Model Details
We choose MICN wang2022micn as the encoder backbone of our model on real-world datasets. Specifically, given the MICN extracts the hidden feature, we apply a variational inference block and then an MLP-based decoder. Architecture details of the proposed method are shown in Table A4.
@@ F.1.1 Data Generation Process
We provide 6 synthetic datasets, with 3,5,8,8,8,16 latent variables from A to F. For dataset A, we have $W_{A}=[[1,1,0],[0,1,0],[0,0,1]]$. For dataset $B$, we have $W_{B}$ as an eye matrix with 2 extra nonzero entries. For dataset $C,F$, we have $W_{C},W_{F}$ as eye matrices. For datasets $D$ and $E$, $W_{D}$ and $W_{E}$ are dense matrices with all nonzero entries, which are generated by the data generator of TDRL(yao2022temporally). When it comes to $V$, we have and only have $V_{i-1,i}=1\ \ \forall i>0$ for dataset $A,B,C,E,F$. We also set $V_{D}=\mathbf{0}$, which means that there are no instantaneous effects.
@@ H.1 Analysis of High-Dimensional Data
Identifying causal relationships in high-dimensional latent variable settings is a well-recognized challenge in the causal inference community (lopez2022large; cheng2024cuts+). As the dimensionality increases, the complexity of identifying causal structures grows due to the expanding search space. We conduct experiments on both synthetic and real-world datasets to verify it and propose a potential solution.
@@ H.2 Noisy Environments
In some real-world scenarios, it is possible for the mixing process not to be invertible, for example, if the mixing process is highly noisy in each observed process, which might be the case in financial data (hu2008instrumental). There are some developments relying on the additive noise assumptions. If one makes strong assumptions on the noise, such as assuming an additive noise model, it is possible to develop a certain type of identifiability just like the extension of nonlinear ICA to the additive noise model case in (khemakhem2020variational). However, general approaches to deal with non-parametric noise terms to developed in this field in the future. To further evaluate the insight of this potential solution, we have conducted experiments on synthetic data. Specifically, we first follow Equation (2) to generate latent variables with temporal and instantaneous dependencies and generate observed variables by ${\mathbf{x}}_{t}={\mathbf{g}}({\mathbf{z}}_{t},\varepsilon)$, where $\varepsilon$ is the noise introduced to the mixing process. Experiment results in Table A20 show that our method can still achieve relatively good identifiability results even in a noisy environment, proving the potential of our insight.
@@ H.4 Subsampled Time Series
Sequentially, we consider another case where the time series data are sampled with low resolutions. When the time series data is sampled with low frequency, additional edges are introduced into the Markov network, making it denser. We can further assume a sparse mixing process zheng2022identifiability to achieve identifiability. Specifically, When conditioned on historical information that provides sufficient changes, the sparse mixing procedure assumption imposes structural constraints on the mapping from estimated to true latent variables. This compensates for potentially insufficient distribution changes, enabling identifiability even when the time series data with low-sample regimes. To further evaluate this insight, we also conduct experiments on synthetic datasets with low-sample regimes. Specifically, we first generate time series data with temporal and instantaneous dependencies. Then we randomly subsample the synthetic time series data. To enforce the sparsity of the estimated mixing process, we add an extra L1 constraint regarding the partial differential between the estimated observed and latent variables, i.e. $|\frac{\partial\hat{{\mathbf{x}}}_{t}}{\partial\hat{{\mathbf{z}}}_{t}}|_{1}$. Experiment results are shown in Table A22.

### v4

@@ B.1 Relationships between Ground-truth and Estimated Latent Variables
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})$ denotes all pairwise indice such that $z_{t,i},z_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $2n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ different values of ${\mathbf{z}}_{t-1,m}$, such that the $2n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ values of vector functions $w(m)$ are linearly independent.
@@ B.2 Extension to Multiple Lags and Sequence Lengths
For the sake of simplicity, we consider only one special case with $\tau=1$ and $L=2$ in Theorem A2. Our identifiability theorem can be actually extended to arbitrary lags and subsequences easily. For any given $\tau$, and subsequence which is centered at ${\mathbf{z}}_{t}$ with previous $lo$ and following $hi$ steps, i.e., ${\mathbf{z}}_{t}=\{{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau}\}$. In this case, the vector function $w(i,j,m)$ in Sufficient Variability Assumption should be modified as
@@ B.3 Identifiability of Latent Variables
First, we demonstrate that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero (Zhang et al., 2024). By contradiction, if the product of the diagonal entry of an invertible matrix $A$ is zero for every row permutation, then we have Equation
@@ Appendix C Discussion of Assumptions
First, The smooth and positive density assumption is standard in the literature on nonlinear ICA (Khemakhem et al., 2020a; Yao et al., 2022; Kong et al., 2022), meaning that the latent variables ${\mathbf{z}}_{t}$ change continuously based on historical information. For instance, in a weather dataset, temperature varies smoothly over time. However, this assumption may be violated if we cannot fully capture the transition probabilities from the observations. To mitigate this, we can sample a larger dataset to better estimate the latent causal process.
Second, the sufficient variability assumption has also been commonly adopted for the existing identifiable results of temporally causal representation learning (Yao et al., 2021; 2022; Chen et al., 2024a). This assumption describes the changeability of latent variables. Take human motion forecasting as an example, the latent variable may represent joint location at different time steps. The linear independence of the latent variables means that the changes of each joint cannot be linearly represented by others. Besides, while the sufficiency assumption is foundational to the theory of identifiability, it is not excessively restrictive. Even in cases where this assumption is not fully satisfied, it is still possible to achieve a degree of subspace identifiability (Kong et al., 2022).
Finally, the sparse latent process is the key assumption of our method, which is common in real-world scenarios. Intuitively, in human motion forecasting, the joints of humans can be considered as latent variables. Since there are few connections among the joints of the human body, the latent process described by the skeleton motion trajectory is sparse. Even though the sparsity assumption is not completely satisfied, we can still attain a subspace level of identifiability (Li et al., 2024). In this case, each true variable can be a function of, at most, an estimated version of its corresponding variable and those within the intimate set. Let us provide a simple example here. In a video of a moving car, it might be hard to have individual identifiability of the separate car wheels and car body; however, they can be considered as essential parts of the macro variable ’car’. This macro representation might be sufficient for the purpose of modeling the interactions between the car and other objects.
@@ E.3 Model Details
We choose MICN Wang et al. (2022) as the encoder backbone of our model on real-world datasets. Specifically, given the MICN extracts the hidden feature, we apply a variational inference block and then an MLP-based decoder. Architecture details of the proposed method are shown in Table A4.
@@ F.1.1 Data Generation Process
We provide 6 synthetic datasets, with 3,5,8,8,8,16 latent variables from A to F. For dataset A, we have $W_{A}=[[1,1,0],[0,1,0],[0,0,1]]$. For dataset $B$, we have $W_{B}$ as an eye matrix with 2 extra nonzero entries. For dataset $C,F$, we have $W_{C},W_{F}$ as eye matrices. For datasets $D$ and $E$, $W_{D}$ and $W_{E}$ are dense matrices with all nonzero entries, which are generated by the data generator of TDRL(Yao et al., 2022). When it comes to $V$, we have and only have $V_{i-1,i}=1\ \ \forall i>0$ for dataset $A,B,C,E,F$. We also set $V_{D}=\mathbf{0}$, which means that there are no instantaneous effects.
@@ H.1 Analysis of High-Dimensional Data
Identifying causal relationships in high-dimensional latent variable settings is a well-recognized challenge in the causal inference community (Lopez et al., 2022; Cheng et al., 2024). As the dimensionality increases, the complexity of identifying causal structures grows due to the expanding search space. We conduct experiments on both synthetic and real-world datasets to verify it and propose a potential solution.
@@ H.2 Noisy Environments
In some real-world scenarios, it is possible for the mixing process not to be invertible, for example, if the mixing process is highly noisy in each observed process, which might be the case in financial data (Hu and Schennach, 2008). There are some developments relying on the additive noise assumptions. If one makes strong assumptions on the noise, such as assuming an additive noise model, it is possible to develop a certain type of identifiability just like the extension of nonlinear ICA to the additive noise model case in (Khemakhem et al., 2020a). However, general approaches to deal with non-parametric noise terms to developed in this field in the future. To further evaluate the insight of this potential solution, we have conducted experiments on synthetic data. Specifically, we first follow Equation (2) to generate latent variables with temporal and instantaneous dependencies and generate observed variables by ${\mathbf{x}}_{t}={\mathbf{g}}({\mathbf{z}}_{t},\varepsilon)$, where $\varepsilon$ is the noise introduced to the mixing process. Experiment results in Table A20 show that our method can still achieve relatively good identifiability results even in a noisy environment, proving the potential of our insight.
@@ H.4 Subsampled Time Series
Sequentially, we consider another case where the time series data are sampled with low resolutions. When the time series data is sampled with low frequency, additional edges are introduced into the Markov network, making it denser. We can further assume a sparse mixing process Zheng et al. (2022) to achieve identifiability. Specifically, When conditioned on historical information that provides sufficient changes, the sparse mixing procedure assumption imposes structural constraints on the mapping from estimated to true latent variables. This compensates for potentially insufficient distribution changes, enabling identifiability even when the time series data with low-sample regimes. To further evaluate this insight, we also conduct experiments on synthetic datasets with low-sample regimes. Specifically, we first generate time series data with temporal and instantaneous dependencies. Then we randomly subsample the synthetic time series data. To enforce the sparsity of the estimated mixing process, we add an extra L1 constraint regarding the partial differential between the estimated observed and latent variables, i.e. $|\frac{\partial\hat{{\mathbf{x}}}_{t}}{\partial\hat{{\mathbf{z}}}_{t}}|_{1}$. Experiment results are shown in Table A22.